# Reducing quantum error correction overhead using soft information

### v1

@@ IV.1.1 Superconducting qubit noise model and simulations
We use the SI1000 noise model from Google Quantum AI [10, 33] as our baseline circuit-level noise model for superconducting qubits. This model is parametrised by the two-qubit gate depolarising error probability $p$, with single-qubit Clifford gates having an error probability $p/10$, initialisation errors in the $Z$-basis having probability $2p$ and $Z$-basis measurement errors having probability $p_{M}=5p$. In our model, we decompose the measurement error probability $p_{M}$ into a contribution $p_{B}$ from a quantum bit-flip error occurring during the measurement process and a contribution with error probability $p_{\textrm{S}}$ from a soft measurement error at the classification stage. We fix $p_{B}=p$ and choose $p_{\textrm{S}}$ based on the noise regime such that the overall measurement error probability satisfies
$p_{M}=p\left(1-p_{\textrm{S}}\right)+p_{\textrm{S}}\left(1-p\right)$ (6)
where in the LF-regime, $p_{\textrm{S}}=p$ leads to $p_{M}\sim 2p$ and in the HF-regime $p_{\textrm{S}}=5p$ leads to $p_{M}\sim 6p$, both taken in the limit $p\ll 1$.
The average classification error probability $p_{\textrm{S}}$ is computed from the overlap integral of the soft measurement response functions $f^{(0)}(\mu)$ and $f^{(1)}(\mu)$. Here we define $f^{(0)}(\mu)$ and $f^{(1)}(\mu)$ according to the dispersive readout model described in Refs. [11, 8]. Given a qubit measurement with measurement time $\tau_{M}$, amplitude damping time $\text{T}_{1}$ and signal-to-noise ratio SNR, the probability density function $f^{(0)}(\mu)$ of the soft outcome $\mu$ for a qubit measured in the ideal outcome $\bar{\mu}=0$ is given by
@@
$\begin{split}f^{(\bar{\mu}=1)}(\mu)&=\sqrt{\frac{\text{SNR}}{4\pi}}\exp\left(-\frac{\text{SNR}}{4}\left(\mu+1\right)^{2}-\frac{\tau_{M}}{\text{T}_{1}}\right)\\ &-\left[g^{(\bar{\mu}=0)}(\mu)-g^{(\bar{\mu}=1)}(\mu)\right]\frac{\tau_{M}}{4\text{T}_{1}}\\ &\times\exp\left(\frac{1}{4\times\text{SNR}}\left(\frac{\tau_{M}}{\text{T}_{1}}\right)^{2}+\frac{\tau_{M}}{2\text{T}_{1}}\left(\mu-1\right)\right)\end{split}$ (8)
@@
$g^{(\bar{\mu}=0)}(\mu)$ $=\text{erfc}\left[\sqrt{\frac{1}{4\times\text{SNR}}}\left(\frac{\tau_{M}}{\text{T}_{1}}\right)+\sqrt{\frac{\text{SNR}}{4}}\left(\mu-1\right)\right]$ (9)
$g^{(\bar{\mu}=1)}(\mu)$ $=\text{erfc}\left[\sqrt{\frac{1}{4\times\text{SNR}}}\left(\frac{\tau_{M}}{\text{T}_{1}}\right)+\sqrt{\frac{\text{SNR}}{4}}\left(\mu+1\right)\right]$ (10)
@@
$\text{SNR}=\frac{\tau_{M}}{2\tau_{F}}$ (11)
where $\tau_{F}$ is a characteristic fluctuation time of the readout signal. In the limit where the measurement time is much shorter than the amplitude damping time $t_{M}\ll\text{T}_{1}$, the classification error probability $p_{\textrm{S}}$ can be easily calculated from
@@
In the time-dependence analysis of section II.3.2, we use the SI1000 model as a starting point for our circuit-level noise model and incorporate the measurement time $\tau_{M}$ into the noise channel probabilities as follows. We maintain the same error channel probabilities for single-qubit gates, two-qubit gates and reset operations as SI1000. Time-dependence is introduced by incorporating idling noise to qubits that are not acted upon by gates, measurements or resets, with the probability of the idling error channel determined by the specific duration of each operation. We set $\tau_{1q}=20~{}\text{ns}$ for single-qubit gates, $\tau_{2q}=40~{}\text{ns}$ for two-qubit gates and $\tau_{R}=40~{}\text{ns}$ for resets. Based on the duration $\tau$ of each operation, single-qubit Pauli noise with components $(P_{X},P_{Y},P_{Z})$ is added to the idling qubits according to the following probabilities, as taken from Ref. [36]:
@@
where $\textrm{T}_{1}$ is the longitudinal relaxation time and $\textrm{T}_{2}$ is the dephasing time. To ensure measurement errors in the time-dependent model match the probability $p_{B}=p$ as seen in the SI1000 model, we parametrise the measurement bit-flip channel probability via
$p_{B}(\tau_{M})=1-\exp\left(-\tau_{M}/\tau_{D}\right)$ (14)
where $\tau_{D}$ is a characteristic depolarising timescale of the device. We choose $\tau_{D}$ such that the measurement bit-flip probability $p_{B}$ in the time-dependent noise model results in an error rate of $p$ given a reference measurement time $\tau_{M}=500~{}\text{ns}$.
For simplicity, we assume that the measurement response of every qubit is identical, i.e. the parameters $\text{T}_{1}$ and $\tau_{F}$ are uniform across all qubits in the device. This is a limitation of our model, and in the real world manufacturing inconsistencies may lead to non-uniform measurement characteristics for different qubits. As a consequence, optimal readout fidelity may be achieved with different readout durations $\tau_{M}$ based on the properties of each qubit. Future investigations may illuminate how non-uniform qubit measurement responses affect the effectiveness of the soft decoding techniques outlined in this work.
@@ IV.1.2 Neutral atom noise model and simulations
$f^{(\bar{\mu}=1)}(\mu;t)$ $=e^{-(\eta R_{0}+R_{bg}+R_{b\to d})t}\frac{(\eta R_{0}+R_{bg})^{\mu}t^{\mu}}{\mu!}+\left(\frac{R_{b\to d}e^{-R_{bg}t}}{\eta R_{0}+R_{b\to d}}\right)\left(\frac{\eta R_{0}}{\eta R_{0}+R_{b\to d}}\right)^{\mu}$
×$\quad\times\left[\sum_{k=0}^{\mu}\frac{(\eta R_{0}+R_{b\to d})^{k}(R_{bg}t)^{k}}{k!(\eta R_{0})^{k}}-e^{-(\eta R_{0}+R_{b\to d})t}\sum_{k=0}^{n}\frac{(\eta R_{0}+R_{b\to d})^{k}(\eta R_{0}+R_{bg})^{k}t^{k}}{k!(\eta R_{0})^{k}}\right]$ (15a)
$f^{(\bar{\mu}=0)}(\mu;t)$ $=e^{-(R_{bg}+R_{d\to b})t}\frac{(R_{bg}t)^{\mu}}{\mu!}+\left(\frac{R_{d\to b}e^{-R_{bg}t}}{\eta R_{0}-R_{d\to b}}\right)\left(\frac{\eta R_{0}}{\eta R_{0}-R_{d\to b}}\right)^{\mu}$
×$\quad\times\left[e^{-R_{d\to b}t}\sum_{k=0}^{\mu}\frac{(\eta R_{0}-R_{d\to b})^{k}(R_{bg}t)^{k}}{k!(\eta R_{0})^{k}}-e^{-\eta R_{0}t}\sum_{k=0}^{\mu}\frac{(\eta R_{0}-R_{d\to b})^{k}(\eta R_{0}+R_{bg})^{k}t^{k}}{k!(\eta R_{0})^{k}}\right]$ (15b)
The main parameters for this readout model include the scattering rate of an atom in the bright state $R_{0}$, the detection efficiency $\eta$, the background scattering rate $R_{\textrm{bg}}$, the scattering rate when atoms transition from a bright to a dark state $R_{\textrm{bd}}$, and the rate of scattering when atoms transition from a dark to a bright state $R_{\textrm{db}}$. In a setting where background scattering and transition errors are low, the longer the readout pulse the smaller the overlap is between the two distributions, resulting in a lower soft measurement flip probability. The distributions for the bright and the dark states are Poissonian, with an effective mean no. of photons are $\lambda_{\textrm{bright}}=(\eta R_{0}+R_{\textrm{bg}})\tau_{M}$ in a bright state and $\lambda_{\textrm{dark}}=R_{\textrm{bg}}\tau_{M}$ in a dark state. To reflect a realistic neutral atom readout chain [23], our simulations are configured with $\eta=0.1\%$, $\tau_{M}=100~{}\mu\textrm{s}$, $R_{0}=10^{7}$, $R_{\textrm{bg}}=10^{3}$, $R_{\textrm{bd}}=960$ and $R_{\textrm{db}}=2$. The qubits in neutral atoms are naturally identical and have long coherence times, therefore, it is safe to assume uniform measurement response functions across all qubits in a simulation.
@@ IV.2 Estimating the error suppression rate
$p_{L}=\frac{1}{2}\left[(1-(1-2\epsilon_{L})^{T}\right]$ (16)
where $\epsilon_{L}$ is the per-round logical error rate [10]. Solving the above for $\epsilon_{L}$ in the limit of low error rates $\epsilon_{L}\ll 1$ as done in Ref. [15] gives the approximate relation
$\epsilon_{L}\approx\frac{p_{L}}{T}\quad.$ (17)
To extract $\Lambda$ as defined in Eq. 4, we first compute the per-round error rate $\epsilon_{L}(d)$ from Eq. 17 for varying code distances $d$, and take a least-squares fit of $x=(d+1)/2$ against $y=\log\left(\epsilon_{L}(d)\right)$. The resulting fit takes the form
$\log\left[\epsilon_{L}(d)\right]=-\log(\Lambda)\times d+C$ (18)
where $C$ is a constant offset, and we can extract $\Lambda$ from the coefficient in front of $d$. To estimate the uncertainty in $\Lambda$, we use the Sinter library (a sub-package of Stim [32]) to evaluate hypotheses with a likelihood within a factor of 10 of the maximum likelihood hypothesis, given the sampled data. This method gives us error bars on the logical error probability $p_{L}$, avoiding the pitfall of making overconfident estimates about the accuracy of the logical error probability at low error rates. We propagate the Sinter-derived uncertainty $\Delta p_{L}$ to $\epsilon_{L}$ via Eq. 17 and use the resulting error bars in the curve fitting process of Eq. 18 to determine the standard error in $\Lambda$.
@@ IV.4 Soft Belief Propagation Decoder
For our purposes, we focus on the Calderbank-Shor-Steane (CSS) subclass of qLDPC codes under a circuit-level noise model. We show an example of the Tanner graph for a distance-$3$ surface code in Fig. 7. The Tanner graph for a circuit-level noise model can be represented as $\mathcal{T}=(Q,S,E)$ containing a set of variables nodes ($Q_{i}$) which store the information about the error mechanisms based on the noise model used, while the check nodes ($S_{i}$) correspond to the detectors and $E$ is the edge set. If the error mechanism flips the detectors there exists an edge between $Q_{i}$ and $S_{i}$. In the case of multiple error mechanisms flipping the same detectors, we merge these error mechanisms into one variable node—i.e merge the equivalent variable nodes. The BP decoder is initialized with log-likelihood ratios, also known as priors, the probabilities of which we obtain from the circuit-level noise model. The use of log-likelihood ratios is to ensure numerical stability of BP.
@@ VII Author contributions
J. M led the research and writing of the manuscript, modelled soft information in the context of superconducting qubits, incorporated soft incorporation decoding into the local clustering decoder and developed software to simulate and analyse decoding performance under varying noise regimes. E.S.M led the implementation of soft-information-aware belief propagation algorithm, designed and performed simulations of QEC experiments with the soft belief propagation decoder and modelled QEC experiments on neutral atom platforms. The manuscript was jointly written by both authors.

### v2

@@ IV.1.1 Superconducting qubit noise model and simulations
We use the SI1000 noise model from Google Quantum AI [10, 34] as our baseline circuit-level noise model for superconducting qubits. This model is parametrised by the two-qubit gate depolarising error probability $p$, with single-qubit Clifford gates having an error probability $p/10$, initialisation errors in the $Z$-basis having probability $2p$ and $Z$-basis measurement errors having probability $p_{\textrm{M}}=5p$. In our model, we decompose the measurement error probability $p_{\textrm{M}}$ into a contribution $p_{\textrm{B}}$ from a quantum bit-flip error occurring during the measurement process and a contribution with error probability $p_{\textrm{S}}$ from a soft measurement error at the classification stage. We fix $p_{\textrm{B}}=p$ and choose $p_{\textrm{S}}$ based on the noise regime such that the overall measurement error probability satisfies
$p_{\textrm{M}}=p\left(1-p_{\textrm{S}}\right)+p_{\textrm{S}}\left(1-p\right)$ (6)
where in the LF-regime, $p_{\textrm{S}}=p$ leads to $p_{\textrm{M}}\sim 2p$ and in the HF-regime $p_{\textrm{S}}=5p$ leads to $p_{\textrm{M}}\sim 6p$, both taken in the limit $p\ll 1$.
The average classification error probability $p_{\textrm{S}}$ is computed from the overlap integral of the soft measurement response functions $f^{(0)}(\mu)$ and $f^{(1)}(\mu)$. Here we define $f^{(0)}(\mu)$ and $f^{(1)}(\mu)$ according to the dispersive readout model described in Refs. [12, 8]. Given a qubit measurement with measurement time $\tau_{\textrm{M}}$, amplitude damping time $\text{T}_{1}$ and signal-to-noise ratio SNR, the probability density function $f^{(0)}(\mu)$ of the soft outcome $\mu$ for a qubit measured in the ideal outcome $\bar{\mu}=0$ is given by
@@
$\begin{split}f^{(\bar{\mu}=1)}(\mu)&=\sqrt{\frac{\text{SNR}}{4\pi}}\exp\left(-\frac{\text{SNR}}{4}\left(\mu+1\right)^{2}-\frac{\tau_{\textrm{M}}}{\text{T}_{1}}\right)\\ &-\left[g^{(\bar{\mu}=0)}(\mu)-g^{(\bar{\mu}=1)}(\mu)\right]\frac{\tau_{\textrm{M}}}{4\text{T}_{1}}\\ &\times\exp\left(\frac{1}{4\times\text{SNR}}\left(\frac{\tau_{\textrm{M}}}{\text{T}_{1}}\right)^{2}+\frac{\tau_{\textrm{M}}}{2\text{T}_{1}}\left(\mu-1\right)\right)\end{split}$ (8)
@@
$g^{(\bar{\mu}=0)}(\mu)$ $=\text{erfc}\left[\sqrt{\frac{1}{4\times\text{SNR}}}\left(\frac{\tau_{\textrm{M}}}{\text{T}_{1}}\right)+\sqrt{\frac{\text{SNR}}{4}}\left(\mu-1\right)\right]$ (9)
$g^{(\bar{\mu}=1)}(\mu)$ $=\text{erfc}\left[\sqrt{\frac{1}{4\times\text{SNR}}}\left(\frac{\tau_{\textrm{M}}}{\text{T}_{1}}\right)+\sqrt{\frac{\text{SNR}}{4}}\left(\mu+1\right)\right]$ (10)
@@
$\text{SNR}=\frac{\tau_{\textrm{M}}}{2\tau_{\textrm{F}}}$ (11)
where $\tau_{\textrm{F}}$ is a characteristic fluctuation time of the readout signal. In the limit where the measurement time is much shorter than the amplitude damping time $\tau_{\textrm{M}}\ll\text{T}_{1}$, the classification error probability $p_{\textrm{S}}$ can be easily calculated from
@@
In the time-dependence analysis of section II.3.2, we use the SI1000 model as a starting point for our circuit-level noise model and incorporate the measurement time $\tau_{\textrm{M}}$ into the noise channel probabilities as follows. We maintain the same error channel probabilities for single-qubit gates, two-qubit gates and reset operations as SI1000. Time-dependence is introduced by incorporating idling noise to qubits that are not acted upon by gates, measurements or resets, with the probability of the idling error channel determined by the specific duration of each operation. We set $\tau_{\textrm{1q}}=20~{}\text{ns}$ for single-qubit gates, $\tau_{\textrm{2q}}=40~{}\text{ns}$ for two-qubit gates and $\tau_{\textrm{R}}=40~{}\text{ns}$ for resets. Based on the duration $\tau$ of each operation, single-qubit Pauli noise with components $(P_{X},P_{Y},P_{Z})$ is added to the idling qubits according to the following probabilities, as taken from Ref. [37]:
@@
where $\textrm{T}_{1}$ is the longitudinal relaxation time and $\textrm{T}_{2}$ is the dephasing time. To ensure measurement errors in the time-dependent model match the probability $p_{\textrm{B}}=p$ as seen in the SI1000 model, we parametrise the measurement bit-flip channel probability via
$p_{\textrm{B}}(\tau_{\textrm{M}})=1-\exp\left(-\tau_{\textrm{M}}/\tau_{\textrm{D}}\right)$ (14)
where $\tau_{\textrm{D}}$ is a characteristic depolarising timescale of the device. We choose $\tau_{\textrm{D}}$ such that the measurement bit-flip probability $p_{B}$ in the time-dependent noise model results in an error rate of $p$ given a reference measurement time $\tau_{\textrm{M}}=500~{}\text{ns}$.
For simplicity, we assume that the measurement response of every qubit is identical, i.e. the parameters $\text{T}_{1}$ and $\tau_{\textrm{F}}$ are uniform across all qubits in the device. This is a limitation of our model, and in the real world manufacturing inconsistencies may lead to non-uniform measurement characteristics for different qubits. As a consequence, optimal readout fidelity may be achieved with different readout durations $\tau_{\textrm{M}}$ based on the properties of each qubit. Future investigations may illuminate how non-uniform qubit measurement responses affect the effectiveness of the soft decoding techniques outlined in this work.
@@ IV.1.2 Neutral atom noise model and simulations
$f^{(\bar{\mu}=1)}(\mu;t)$ $=e^{-(\eta R_{0}+R_{\textrm{bg}}+R_{b\to d})t}\frac{(\eta R_{0}+R_{\textrm{bg}})^{\mu}t^{\mu}}{\mu!}+\left(\frac{R_{b\to d}e^{-R_{\textrm{bg}}t}}{\eta R_{0}+R_{b\to d}}\right)\left(\frac{\eta R_{0}}{\eta R_{0}+R_{b\to d}}\right)^{\mu}$
×$\quad\times\left[\sum_{k=0}^{\mu}\frac{(\eta R_{0}+R_{b\to d})^{k}(R_{\textrm{bg}}t)^{k}}{k!(\eta R_{0})^{k}}-e^{-(\eta R_{0}+R_{b\to d})t}\sum_{k=0}^{n}\frac{(\eta R_{0}+R_{b\to d})^{k}(\eta R_{0}+R_{\textrm{bg}})^{k}t^{k}}{k!(\eta R_{0})^{k}}\right]$ (15a)
$f^{(\bar{\mu}=0)}(\mu;t)$ $=e^{-(R_{\textrm{bg}}+R_{d\to b})t}\frac{(R_{\textrm{bg}}t)^{\mu}}{\mu!}+\left(\frac{R_{d\to b}e^{-R_{\textrm{bg}}t}}{\eta R_{0}-R_{d\to b}}\right)\left(\frac{\eta R_{0}}{\eta R_{0}-R_{d\to b}}\right)^{\mu}$
×$\quad\times\left[e^{-R_{d\to b}t}\sum_{k=0}^{\mu}\frac{(\eta R_{0}-R_{d\to b})^{k}(R_{\textrm{bg}}t)^{k}}{k!(\eta R_{0})^{k}}-e^{-\eta R_{0}t}\sum_{k=0}^{\mu}\frac{(\eta R_{0}-R_{d\to b})^{k}(\eta R_{0}+R_{\textrm{bg}})^{k}t^{k}}{k!(\eta R_{0})^{k}}\right]$ (15b)
The main parameters for this readout model include the scattering rate of an atom in the bright state $R_{0}$, the detection efficiency $\eta$, the background scattering rate $R_{\textrm{bg}}$, the scattering rate when atoms transition from a bright to a dark state $R_{b\to d}$, and the rate of scattering when atoms transition from a dark to a bright state $R_{d\to b}$. In a setting where background scattering and transition errors are low, the longer the readout pulse the smaller the overlap is between the two distributions, resulting in a lower soft measurement flip probability. The distributions for the bright and the dark states are Poissonian, with an effective mean no. of photons are $\lambda_{\textrm{bright}}=(\eta R_{0}+R_{\textrm{bg}})\tau_{\textrm{M}}$ in a bright state and $\lambda_{\textrm{dark}}=R_{\textrm{bg}}\tau_{\textrm{M}}$ in a dark state. To reflect a realistic neutral atom readout chain [24], our simulations are configured with $\eta=0.1\%$, $\tau_{\textrm{M}}=100~{}\mu\textrm{s}$, $R_{0}=10^{7}$, $R_{\textrm{bg}}=10^{3}$, $R_{b\to d}=960$ and $R_{d\to b}=2$. The qubits in neutral atoms are naturally identical and have long coherence times, therefore, it is safe to assume uniform measurement response functions across all qubits in a simulation.
@@ IV.2 Estimating the error suppression rate
$P_{\textrm{L}}=\frac{1}{2}\left[(1-(1-2\epsilon_{\textrm{L}})^{T}\right]$ (16)
where $\epsilon_{\textrm{L}}$ is the per-round logical error rate [10]. Solving the above for $\epsilon_{\textrm{L}}$ in the limit of low error rates $\epsilon_{\textrm{L}}\ll 1$ as done in Ref. [16] gives the approximate relation
$\epsilon_{\textrm{L}}\approx\frac{P_{\textrm{L}}}{T}\quad.$ (17)
To extract $\Lambda$ as defined in Eq. 4, we first compute the per-round error rate $\epsilon_{\textrm{L}}(d)$ from Eq. 17 for varying code distances $d$, and take a least-squares fit of $x=(d+1)/2$ against $y=\log\left(\epsilon_{\textrm{L}}(d)\right)$. The resulting fit takes the form
$\log\left[\epsilon_{\textrm{L}}(d)\right]=-\log(\Lambda)\times d+C$ (18)
where $C$ is a constant offset, and we can extract $\Lambda$ from the coefficient in front of $d$. To estimate the uncertainty in $\Lambda$, we use the Sinter library (a sub-package of Stim [33]) to evaluate hypotheses with a likelihood within a factor of 10 of the maximum likelihood hypothesis, given the sampled data. This method gives us error bars on the logical error probability $P_{\textrm{L}}$, avoiding the pitfall of making overconfident estimates about the accuracy of the logical error probability at low error rates. We propagate the Sinter-derived uncertainty $\Delta P_{\textrm{L}}$ to $\epsilon_{\textrm{L}}$ via Eq. 17 and use the resulting error bars in the curve fitting process of Eq. 18 to determine the standard error in $\Lambda$.
@@ IV.4 Soft Belief Propagation Decoder
For our purposes, we focus on the Calderbank-Shor-Steane (CSS) subclass of qLDPC codes under a circuit-level noise model. We show an example of the Tanner graph for a distance-$3$ surface code in Fig. 7. The Tanner graph for a circuit-level noise model can be represented as $\mathcal{T}=(Q,S,E)$ containing a set of variables nodes ($Q_{i}$) which store the information about the error mechanisms based on the noise model used, while the check nodes ($S_{i}$) correspond to the detectors and $E$ is the edge set. If the error mechanism flips the detectors there exists an edge between $Q_{i}$ and $S_{i}$. In the case of multiple error mechanisms flipping the same detectors, we merge these error mechanisms into one variable node—i.e. merge the equivalent variable nodes. The BP decoder is initialized with log-likelihood ratios, also known as priors, the probabilities of which we obtain from the circuit-level noise model. The use of log-likelihood ratios is to ensure numerical stability of BP.
@@ VII Author contributions
JM led the research and writing of the manuscript, modelled soft information in the context of superconducting qubits, incorporated soft incorporation decoding into the local clustering decoder and developed software to simulate and analyse decoding performance under varying noise regimes. ESM led the implementation of soft-information-aware belief propagation algorithm, designed and performed simulations of QEC experiments with the soft belief propagation decoder and modelled QEC experiments on neutral atom platforms. The manuscript was jointly written by both authors.